Use the data in the table below to complete parts​ (a) through​ (c).

x

5

7

10

11

15

18

18

40

y

33

36

33

30

26

25

22

2

A.) Construct a scatterplot of the data

B.) Identify any possible outliners

C.

c) Determine if the point is influential. The change in slope or intercept is significant if it is larger than​ 10%.

The point

is not

is

an influential point because the slopes with the point included and without the point included

are

are not

significantly​ different, and the intercepts

are

are not

significantly different.
Answer #1
A) Scatter Plot

B) Point X=40 and Y=2 is a possible outlier

C) Find the regression slope using Excel Data Analysis Regression tool
with and without the point
Slope with the point included = -0.964
Intercept with the point included = 40.816
Slope without the point included = -0.919
Intercept with the point included = 40.311

Difference in slopes = 0.964-0.919 = 0.045
Difference in slopes is 4.67% which is less than 10%
Difference in intercepts = 40.816 - 40.311 = 0.505
Difference in intercepts is 1.24% which is less than 10%

The point is not an influential point because
the slopes with the point included and without the point included
are not significantly different
and the intercepts are not significantly different